Suppose you purchased 19 shares of Disney stock for ​$23.95 per share on May​ 1, 2012. On September 1 of the same​ year, you sold 19 shares of the stock for ​$22.00. Calculate the​ holding-period dollar gain for the shares you​ sold, assuming no dividend was​ distributed, and the​ holding-period rate of return.
Answer #1

Solution:

Holding period dollar gain = Number of shares(Selling price - Buying Price)

= 19 (23.95 - 22.00)

= $37.05

Rate of return = ((23.95 - 22.00)/22.00)*100 = 8.86%

Holding period rate of return = 8.86 (4/12) = 2.95%